OpenStudy (anonymous):

What is the value of the variable that makes the statement true? a + (–7) + 3 = 0 4 10 –10 –4

OpenStudy (anonymous):

c

OpenStudy (anonymous):

What is the value of the expression |x + y| when x = –8 and y = –13? –5 –21 5 21

OpenStudy (trojanpoem):

Solve for a: a -4 = 0 a = 4

OpenStudy (trojanpoem):

Plug x, y in the equation: |-8 -13| = |-21| = 21 (abs remove the negative sign)

OpenStudy (anonymous):

@trojanpoem Which expression has a value of 23? –3 – 20 7 – (–16) 10 – 13 –18 – (–5)

OpenStudy (trojanpoem):

1) -3 -20 = -23 2) 7 + 16 = 23 3) -3 4) -23

OpenStudy (anonymous):

so whats the answer

OpenStudy (trojanpoem):

7 - (-16)

OpenStudy (trojanpoem):

I hate giving direct answers, as that never helps :/

OpenStudy (anonymous):

Subtract. –21 – (–13) 8 –8 –34 34

OpenStudy (texaschic101):

-21 -(-13) = -21 + 13 = -8

OpenStudy (anonymous):

What is the value of the expression 1 – r – a when r = 20 and a = 6? –13 –25 –15 –27

OpenStudy (trojanpoem):

1- 20 - 6 = -25

OpenStudy (trojanpoem):

Just substitute the values given for the variables a, b

OpenStudy (anonymous):

What is the value of the expression 5 – c – d when c = 8 and d = –13? 0 –16 –26 10

OpenStudy (trojanpoem):

5 - 8 + 13 = 10

OpenStudy (anonymous):

What is the value of the expression 3 – a – b when a = 15 and b = –3? –15 –21 –9 21

OpenStudy (trojanpoem):

I am not helping that way , are i ?

OpenStudy (anonymous):

no i understand now thanks

OpenStudy (anonymous):

What is the value of the expression 3 – a – b when a = 15 and b = –3? –15 –21 –9 21

OpenStudy (trojanpoem):

3 - 15 + 3 = -9